Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

solved correctly

Expert replies
by sanju09 » Sat Feb 13, 2010 1:12 am
A GMAT’s quant question is given to 3 students who have 1/3, ¼ and ½ chances of solving it correctly. What are the chances that the problem will be solved correctly?
(A) ¼
(B) 1/3
(C) 2/3
(D) ¾
(E) 31/40
The mind is everything. What you think you become. –Lord Buddha



Sanjeev K Saxena
Quantitative Instructor
The Princeton Review - Manya Abroad
Lucknow-226001

www.manyagroup.com
Join the discussion
Source: — Problem Solving |

by fibbonnaci » Sat Feb 13, 2010 1:36 am
the question in essence asks us to find the probability of atleast one solving the question correctly.

atleast questions are best answered using the formula:

1- none of them solve the question.

for none to solve the question:it is the prob of 1st person not solving * 2nd person not solving *3rd person not solving.

probability of 1st person not solving = 2/3
2nd person not solving= 3/4
3rd person not solving = 1/2

2/3*3/4*1/2 =>1/4.

1-1/4 => 3/4 => D
Join the discussion

by ajith » Sat Feb 13, 2010 7:07 am
sanju09 wrote:A GMAT’s quant question is given to 3 students who have 1/3, ¼ and ½ chances of solving it correctly. What are the chances that the problem will be solved correctly?
(A) ¼
(B) 1/3
(C) 2/3
(D) ¾
(E) 31/40
The probability that each of them solve it wrong = 2/3 , 3/4 and 1/2
Since these are independent events chances of these happening together = 2/3*3/4*1/2 = 1/4

Since the probability that all of them getting it wrong is 1/4, probability that any of them getting it correct is 1-1/4 = 3/4
Always borrow money from a pessimist, he doesn't expect to be paid back.
Join the discussion